OpenStudy (anonymous):

I do NOT want the flat out answer , i want help on how to solve the problem . For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. 0.08 0.496 2.48 5

OpenStudy (anonymous):

You can graph at x=3, and x=8 on a graphical calculator and connect the two coordinates and calculate the slope of the line.

OpenStudy (anonymous):

slope=rise/run

OpenStudy (anonymous):

https://www.desmos.com

OpenStudy (anonymous):

it has an undefined slope

OpenStudy (anonymous):

Perhaps you entered the x as something other than exponent?

OpenStudy (anonymous):

Every function gets a slope. You must enter the x as ^x

OpenStudy (anonymous):

oh okay give me a min

OpenStudy (loser66):

f(8) =? f(3) =?

OpenStudy (loser66):

and then \(\dfrac{f(8)-f(3)}{8-3}=?\) That is your answer.

OpenStudy (anonymous):

5 ?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

\(f(x) = 0.01*2^x\) right?